OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? 2x-y=-5 x+3y=22 A. (1, 7) B. (4, 9) C. (7, 19) D. (8, 21)

OpenStudy (anonymous):

@TheSavior

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

A

OpenStudy (anonymous):

how

OpenStudy (anonymous):

check these yourself, 2(1)-7=-5 and 1+ 3(7) = 22

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

wait so I do 2*1-7 and 1+3*7

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and see if they match what it says it equals... if it doesnt it is incorrect.

OpenStudy (anonymous):

I got -5 and 28

OpenStudy (anonymous):

how did you do the second one?

OpenStudy (anonymous):

-5 is right, 28 is not

OpenStudy (anonymous):

1+3*7

OpenStudy (anonymous):

if you do 7*3 + 1 then its 22 but if you do it 1+3*7 then its 28

OpenStudy (anonymous):

no, do PEMDAS, always multiply first

OpenStudy (anonymous):

therefore, 1+3*7= 1+21=22

OpenStudy (anonymous):

I'm so dumb lol and PEMDAS was my favorite thing to do

OpenStudy (anonymous):

could you help me with 3 more

OpenStudy (anonymous):

@TheSavior